last year total number of students in a school was 5000 this year the number of boys increased by 5% and that of girls increased by 3% the total number of students increased by 202 to how many boys and girls were there in a school last year
Answer by ankor@dixie-net.com(22740)

last year total number of students in a school was 5000.
b + g = 5000
g = -b + 5000
this year the number of boys increased by 5% and that of girls increased by 3%.
the total number of students increased by 202 to.
.05b + .03g = 202
how many boys and girls were there in a school last year.
Replace g = -b+5000
.05b + .03(-b+5000) = 202
.05b - .03b + 150 = 202
.02b = 202 - 150
.02b = 52
b = 52/.02
b = 2600 boys
and
b = 5000 - 2600
b = 2400 girls
:
:
Check
.05(2600) = 130
.03(2400) = 72
-------------------
total increase: 202
: